# DPImageBench: A Unified Benchmark for Differentially Private Image Synthesis

### v2

@@ 5.2.3 Extensions to New Algorithms
To add new algorithms, users simply need to implement the abstract base class with three interfaces: (1) pertaining the synthesizer with a public image dataset as input, (2) training the synthesizer with a sensitive image dataset as input, and (3) generating synthetic images. Users can flexibly design their methods inside the interfaces.
To refine existing methods, since DPImageBench isolates the parameter configurations from the algorithm implementation, users typically only need to modify the configuration files to adjust hyperparameters and refine algorithms.
@@ 5.3 Evaluation
5.3.1 Previous Synthesis Evaluations
Table III summarizes the evaluations of the investigated algorithms. We introduce them as follows.
Fidelity. The Fréchet Inception Distance (FID) [23] evaluates the quality and diversity of generated images by measuring the distance between the feature distributions of sensitive and synthetic images within the feature space of a pretrained network. It is used by all papers. Inception Score (IS) [27] evaluates the generated images by assessing how confidently the images are classified and how diverse they are across different classes (mode coverage) [61]. Note that IS only makes sense for natural images (e.g., the ones in the ImageNet dataset [59]).
Utility. Utility considers how synthetic images perform in downstream analytics. The most widely explored task is classification. In particular, current studies train classifiers on synthetic images and evaluate the classification accuracy (Acc) using sensitive test datasets.
Empirical DP testing. Current DP image synthesis methods [22, 4, 5] propose to empirically estimate the privacy leakage from the synthetic images under DP constraints. In PrivImage, authors leverage MIAs [44] to infer the sensitive images for the synthetic images. Lin et al. [5] compared the distribution of distances between synthetic images and their nearest sensitive images to confirm that the synthetic images are not copies of private images.
5.3.2 Fidelity Evaluations
As discussed above, most DP image synthesis methods evaluate the fidelity of synthetic images using the FID [23] and IS [27]. They have certain limitations, such as being insensitive to whether synthetic images memorize training images and lacking alignment with human perception [61]. We establish evaluation from four perspectives.
DPImageBench introduces precision and recall [62], which provide separate scores to offer a clearer understanding of whether a model excels at generating high-quality images (precision) or diverse images (recall). Moreover, DPImageBench also incorporates Feature Likelihood Divergence (FLD) [61], which evaluates the novelty of the synthetic dataset. From a human-level perspective, we introduce additional evaluation metrics ImageReward (IR) [25]. We briefly outline the selected metrics as follows:
FID: FID [23] first uses Inception-v3 [63] to extract features from synthetic and sensitive images. It estimates multivariate Gaussian distributions from these features and calculates the FID score by measuring the Fréchet distance [23] between the two distributions. A lower FID indicates that the generated images are more similar to the real dataset.
IS: The IS uses the Inception v3 [63] network to classify each generated image into predefined classes. A higher IS indicates that the synthetic images are both realistic (the model classifies them with high confidence) and diverse (the images span various classes).
Precision and Recall: We begin by extracting features from both sensitive and synthetic images using the Inception v3. Next, for each synthetic image, we identify the closest sensitive image in the feature space. If the distance between the synthetic image and the sensitive image falls below a specified threshold, we classify it as a “true positive.” Precision is then calculated as . Additionally, for recall, we locate the nearest synthetic image for each sensitive image. Here, a “True Positive” is defined as a pair where the distance between the images is below the threshold, and Recall is calculated as . Achieving higher Precision and Recall is preferable, although there is often a trade-off between them [62].
FLD: We refer to the repository released by Jiralerspong et al. [61]. A lower FLD value is better, with negative values indicating the synthetic dataset outperforms the original.
@@ 5.3.2 Fidelity Evaluations
5.3.3 Utility Evaluations
When evaluating synthetic image utility, many studies use simple classifiers, limiting classification performance and reducing evaluation credibility. For instance, DP-MERF [24] reports the best accuracy of 87.9% on real FashionMNIST dataset, which is lower than the 94.4% achieved by Inception v3 [63]. DPImageBench trains classifiers using methods that achieve accuracy comparable to SOTA methods. We also provide options for the other classification algorithms.
MNIST [64] | 70,000 | 28281 | 10 | 55,000 | 5,000 | 10,000 | Classification | Sensitive |
FashionMNIST [65] | 70,000 | 28281 | 10 | 55,000 | 5,000 | 10,000 | Classification | |
CIFAR-10 [66] | 60,000 | 32323 | 10 | 45,000 | 5,000 | 10,000 | Classification | |
CIFAR-100 [66] | 60,000 | 32323 | 100 | 45,000 | 5,000 | 10,000 | Classification | |
EuroSAT [67] | 27,000 | 64643 | 10 | 21,000 | 2,000 | 4,000 | Classification | |
CelebA [43] | 182,732 | 1281283 | 2 | 145,064 | 17,706 | 19,962 | Classification | |
Camelyon [42] | 337,340 | 64643 | 2 | 269,538 | 32,898 | 34,904 | Classification | |
ImageNet ISLVRC2012 [59] | 1,431,167 | 2562563 | 1,000 | 1,281,167 | 50,000 | 100,000 | Pretraining | Public |
Places365 [68] | 2,141,660 | 2562563 | 365 | 1,803,460 | 17,800 | 320,400 | Pretraining |
5.3.4 Promoting Utility Evaluations
Current utility evaluation methods exhibit drawbacks. As described in Table IV, previous works consider three explicit ways to evaluate utility: (1) Test on a fixed epoch (TestFix): Train the classifier up to a predetermined epoch, then directly evaluate it on the sensitive test set. (2) Sensitive test set as val (SenV): Train classifiers on the synthetic images, and use test sets without separate validation sets for tuning classifier to select optimal parameters. (3) Synthetic images as val (SynV): Use the synthetic images as the validation set for classifier selection.
@@ 5.3.4 Promoting Utility Evaluations
As presented in Table VI, previous works both have drawbacks as follows,
@@
SenV: Even worse, as Figure 3 illustrates, ‘SenV’ leads to two key issues: (1) Bias in Test Accuracy: Using the Test set for model selection biases results in higher accuracy. (2) Privacy Risks: Model selection based on test or validation metrics from sensitive datasets may breach DP guarantees [69], as the test and validation sets share distributions with the training data.
SynV: Due to the distribution gap between synthetic images and sensitive images, using synthetic images as the validation set to select classifiers for the sensitive set may underestimate the utility.
DPImageBench presents the solutions to overcome the three issues presented in Table VI as follows.
To address the tuning and the bias issues, Figure 3 illustrates that we use the sensitive validation set (as detailed in Table V), which is distinct from the sensitive training and test set, for classifier tuning. However, the validation set shares the same distribution as the sensitive training data, potentially still violating DP [69].
To satisfy DP, we use the “report noisy max” method [70]. We select the best classifier checkpoint that has the most correct samples. To satisfy -DP, we add the Laplace noise from to their correct number and select the best classifier checkpoint based on the noisy numbers. The best classifier is used to calculate the test set accuracy. Since the training and validation sets are two disjoint sets, according to the parallel composition theorem [71], the ultimate privacy guarantee depends only on the worst of the guarantees of each analysis, , where is the privacy costs of the validation set and is the privacy cost of training the synthesizer on the training set. DPImageBench sets without additional privacy cost. Thus, the final privacy cost is still [52]. We abbreviate this method as ‘Noisy SenV’.
In DPImageBench, we provide ‘SynV’ and ‘Noisy SenV’, the two methods that satisfy DP, avoid bias, and support tuning, as the utility evaluation metrics.
5.3.5 Empirical DP Testing
Empirical DP testing is an experimental approach to auditing the privacy properties of an algorithm. This method involves running the algorithm with specially constructed data to empirically assess DP parameters and/or measure the potential exposure of personal information in the input data. The goal is to determine whether the system effectively obscures sensitive details and adheres to the privacy standards of the DP [72].
We chose not to perform empirical DP testing for the following reasons. (1) Lack of suitable methods: Most existing research focuses on auditing DP in classification models [72, 73], models trained on non-image data (e.g., tabular data [74]), or the DP-SGD algorithm [75, 76]. Applying these methods to our problem is not straightforward, as we focus on DP image generation, and some of the approaches we evaluate, such as PE, do not use DP-SGD. (2) Theoretical guarantee: DP synthesis algorithms already have the rigorous theoretical privacy guarantee that the sensitive information leaked from the output data is bounded. If the goal is empirical privacy, we do not even need DP in the first place.
6 Experimental Setup
@@
No DP | 99.7 | 94.5 | 93.1 | 74.6 | 97.5 | 97.7 | 87.7 |
@@
6.1 Investigated Datasets
Referring to previous works [4, 5, 32, 24, 31], we assume that the investigated datasets can be divided into sensitive datasets that require protection and public datasets that can be released publicly without privacy concerns. Table V presents the summary of datasets. It is noticed that certain public datasets, such as LAION [77], raise privacy concerns and copyright issues and, therefore, cannot be selected for use as pretraining datasets [78].
We select seven sensitive datasets: (1) MNIST [64], (2) FashionMNIST [65], (3) CIFAR-10 [66], (4) CIFAR-100 [66], (5) EuroSAT [67], (6) CelebA [43], and (7) Camelyon [42], which are widely used in previous works [4, 5, 18]. ImageNet ISLVRC2012 [59] and Places365 [68] datasets are considered as the public datasets. CelebA [43] includes various attributes, like hair, glasses, and gender. Following previous works [4], we explore gender classification tasks. MNIST and FashionMNIST contain grayscale images, while other datasets are color images. We provide details in Appendix C.
We conduct experiments that focus exclusively on conditional generation tasks. The unconditional generation, where images are unlabeled, can be viewed as a special case of conditional generation where the conditioning information is simply absent or treated as a generic label.
6.2 Implementation of Evaluation Metrics
This section outlines the metrics used to assess synthetic images. We generate 60,000 synthetic images for evaluations for each sensitive dataset. Besides, when computing metrics such as FID, IS, Precision, and Recall, it is essential to utilize a pretrained neural network to extract features or classify synthetic images. In DPImageBench, we use the pretrained Inception v3 [63], downloaded from Github.222https://github.com/mseitzer/pytorch-fid/releases/download/fid_weights/pt_inception-2015-12-05-6726825d.pth We elaborate on the hyper-parameter settings of utility and fidelity evaluation metrics in Appendices B.2.4 and B.2.5. We evaluate the utility of synthetic images for image classification tasks. DPImageBench offers three classification algorithms—ResNet, WideResNet [79], and ResNeXt [80]—trained using synthetic images [81]. Accuracy is evaluated on the sensitive test set as shown in Table V.
@@ 7 EMPIRICAL EVALUATIONS
We provide the computational resources and runtime analysis of studied methods in Table XIX of Appendix D.
7.1 The Quality of Synthetic Images (RQ1)
@@
Experiment Design. To answer RQ1, this section first studies the utility and fidelity of synthetic images generated by the studied methods under privacy budgets of . We also provide visualizations to offer intuitive presentations. The synthesizer sizes for the GAN-based and diffusion-based methods are 5.4M and 3.8M. For GAN-based methods, we use BigGAN [82], which is a widely used GAN model, with a default generator size of 3.8M and a discriminator size of 1.6M. Thus, the generators used in GAN-based and diffusion-based methods are nearly identical. For diffusion-based methods, we follow the model size settings from DPDM [31] and PrivImage [4]. Besides, for latent diffusion-based methods such as DP-LDM and DP-LoRA, as described in Section 2.2, LDMs integrate an autoencoder into the diffusion process. The DMs remain consistent with other diffusion-based methods (the size of 3.8M), while the autoencoder follows the implementation provided in the original repositories of DP-LDM [18] and DP-LoRA [19]. We provide more descriptions of network architecture in Appendix B.2.2.
@@
The high-dimensional sensitive datasets with resolutions larger than are resized to for fair comparison, following previous works [32, 4, 31]. We also tested higher image resolutions for CelebA, including and . It is noticed that without specific explanation, the following utility evaluations of this paper are obtained under the ‘Noisy SenV’ method.
@@
Result Analysis. Tables VII and VIII show the Acc of classifiers trained on the synthetic images. The best score of each column is marked with the gray color box. Table IX reports the fidelity results under . Figure 4 illustrates the examples of synthetic images under . Notably, we leverage different experimental settings compared to the original papers, including changes to the model architecture, failure probability , and evaluation methods. These differences can lead to variations in performance relative to the results reported in the original publications. These results highlight the following key takeaways:
@@
Directly tuning classifiers using the sensitive dataset overestimates the utility of synthetic images. For example, in Table VII, for the DP-Kernel method with the sensitive dataset EuroSAT, using the synthetic dataset as the validation set results in an accuracy decrease of 21.2% and 24.9% under . Tables VII and VIII present the average results for Noisy SenV’ and SynV,’ showing a decrease of 1.33% and 4.48% in accuracy under , and 1.12% and 3.00% under , compared to the ‘SenV’ evaluation method. As increases, the quality of synthetic images improves, leading to reduced utility evaluation errors. These results also indicate that the ‘Noisy Senv’ method achieves better utility evaluation than the ‘SynV.’
Our fidelity and utility evaluation provides a more comprehensive view of the quality of synthetic images. IS focuses on evaluating the practical quality of synthetic images. However, its ability to evaluate unnatural images, such as MNIST, FashionMNIST, and EuroSAT, remains limited [61]. In Table IX, the PE generates images with IS scores of 2.79, 5.37, 14.97, 19.65, 8.16, 3.87, and 4.58 for the seven sensitive datasets; six out of seven datasets achieve the highest IS across all methods. However, despite the high practical quality of these images, they are not of good utility, achieving Acc scores of 32.7%, 57.8%, 75.3%, 24.9%, 36.8%, 74.2%, and 64.9%. Five out of seven of them are lower than the best methods.
For DP-LDM and DP-LoRA for MNIST, the Acc under is 95.5% and 97.1%, but the FID is 99.1 and 95.4. The synthetic images retain essential information for downstream models but may lack realistic textures, colors, and fine details, as these methods prioritize class-discriminative features like shapes over photorealism. In their original implementations, DP-LDM and DP-LoRA resize MNIST images to , whereas the original size of MNIST image is . This resizing results in better FID scores reported in their paper compared to DPImageBench. When resizing to , DP-LDM and DP-LoRA in DPImageBench achieve FID scores of 13.4 and 12.1, similar to what was reported in their papers. Precision and Recall are decoupled from FID, meaning that having one value high and the other low does not necessarily improve the FID for synthetic images. For example, in the case of DPGAN with CIFAR-10, the precision is 0.67, the best among all methods. However, its recall is as low as 0.01, and the FID is as high as 138.7. FLD considers the novelty of synthetic images, and better FLD scores generally indicate that most other fidelity metrics are also favorable. The human perspective metric, IR, is consistently negative due to the resolution of synthetic images being , which is lower than that of text-to-image synthetic images presented in paper [25], which are . We provide more details of fidelity evaluations for the Camelyon in Appendix D.
@@
Current methods still face challenges in high-resolution DP image synthesis. Table X shows the Acc and FID of synthetic CelebA with varying resolutions. As the resolution of synthetic images increases from to , the average accuracy and FID of synthetic images drop from 84.0 and 87.5 to 65.6 and 218.4, respectively. Current methods still struggle with synthesizing high-resolution images. However, PE is the least affected by the increased resolutions in synthetic images, as the synthetic images of PE only depend on the capability of foundational APIs. Other methods, such as PrivImage and DP-LDM, which train synthesizers using DP-SGD, fail in high-resolution image synthesis. High-resolution images contain more complex spatial structures, requiring more precise gradients to capture meaningful patterns. However, DP-SGD injects noise into gradients, which disrupts these updates. As image resolution increases, the number of model parameters and/or the number of required model updates also grows, making it harder for DP-SGD to converge. Note that we leverage the default parameter size of synthesizers, which is 3.8M. A larger size of synthesizers may improve the synthetic performance.
7.2 Model Architecture and Privacy Budget (RQ2)

### v3

@@ 5.2.3. Extensions to New Algorithms
To add new algorithms, users simply need to implement the abstract base class with three interfaces: (1) pretraining the synthesizer with a public image dataset as input, (2) training the synthesizer with a sensitive image dataset as input, and (3) generating synthetic images. Users can flexibly design their methods inside the interfaces. To refine existing methods, since DPImageBench isolates the parameter configurations from the algorithm implementation, users typically only need to modify the configuration files to adjust hyperparameters and refine algorithms.
@@ 5.3. Evaluation
5.3.1. Previous Synthesis Evaluations
Table 2 summarizes the evaluations of the investigated algorithms. We introduce them as follows.
Fidelity. The Fréchet Inception Distance (FID) (Heusel et al., 2017) measures the quality and diversity of generated images by comparing feature distributions of sensitive and synthetic images in a pretrained network’s feature space. It is used by all papers. The Inception Score (IS) (Salimans et al., 2016) assesses classification confidence and class diversity (Jiralerspong et al., 2023), but is relevant only for natural images (e.g., ImageNet (Deng et al., 2009)).
Utility. Utility evaluates synthetic images in downstream tasks, primarily classification, where classifiers trained on synthetic images are tested for accuracy (Acc) on sensitive datasets.
Empirical DP testing. Current DP image synthesis methods (Li et al., 2024; Lin et al., 2024) propose to empirically assess privacy leakage from synthetic images. PrivImage uses MIAs (Matsumoto et al., 2023) to infer sensitive images from synthetic ones, while PE (Lin et al., 2024) analyzes distance distributions between synthetic and nearest sensitive images to ensure they aren’t copies.
5.3.2. Fidelity Evaluations
As discussed above, most DP image synthesis methods evaluate the fidelity of synthetic images using the FID (Heusel et al., 2017) and IS (Salimans et al., 2016). They have certain limitations, such as being insensitive to whether synthetic images memorize training images and lacking alignment with human perception (Jiralerspong et al., 2023). We establish evaluation from four perspectives.
DPImageBench uses precision and recall (Kynkäänniemi et al., 2019), which provide separate scores to offer a clearer understanding of whether a model excels at generating high-quality images (precision) or diverse images (recall). Moreover, DPImageBench also incorporates Feature Likelihood Divergence (FLD) (Jiralerspong et al., 2023), which evaluates the novelty of the synthetic dataset. From a human-level perspective, we introduce additional evaluation metrics ImageReward (IR) (Xu et al., 2024). We elaborate on the selected metrics in Appendix B.2.5.
5.3.3. Utility Evaluations
When evaluating synthetic image utility, many studies use simple classifiers, limiting classification performance and reducing evaluation credibility. For instance, DP-MERF (Wang et al., 2024) reports the best accuracy of 87.9% on real FashionMNIST dataset, which is lower than the 94.4% achieved by Inception v3 (Szegedy et al., 2016). DPImageBench trains classifiers using methods that achieve accuracy comparable to SOTA methods. We also provide options for the other classification algorithms.
5.3.4. Promoting Utility Evaluations
Current utility evaluation methods have limitations, as shown in Table 3. Prior studies use three approaches: (1) Test on a fixed epoch (TestFix):Train the classifier to a fixed epoch and test it on the sensitive set. (2) Sensitive test set as val (SenV): Train classifiers on the synthetic images, and use test sets without separate validation sets for tuning classifier to select optimal parameters. (3) Synthetic images as val (SynV): Use synthetic images as the validation set for classifier selection. Table 4 reveals their drawbacks, detailed below,
@@ 5.3.4. Promoting Utility Evaluations
SenV: Even worse, as Figure 3 illustrates, ‘SenV’ leads to two key issues: (1) Bias in Test Accuracy: Using the Test set for model selection biases results in higher accuracy. (2) Privacy Risks: Model selection based on test or validation metrics from sensitive datasets may breach DP guarantees (Ponomareva et al., 2023), as the test and validation sets share distributions with the training data.
SynV: Due to the distribution gap between synthetic images and sensitive images, using synthetic images as the validation set to select classifiers may underestimate the utility.
@@
| No DP | 99.7 | 94.5 | 93.1 | 74.6 | 97.5 | 97.7 | 87.7 | |||||||
DPImageBench presents the solutions to overcome the three issues presented in Table 4 as follows.
To address the tuning and the bias issues, Figure 3 illustrates that we use the sensitive validation set (as detailed in Table 16), which is distinct from the sensitive training and test set, for classifier tuning. However, the validation set shares the same distribution as the sensitive training data, potentially still violating DP (Ponomareva et al., 2023).
To satisfy DP, we use the “report noisy max” method (Tang et al., [n. d.]). We select the best classifier checkpoint that has the most correct samples. To satisfy -DP, we add the Laplace noise from to their correct number and select the best classifier checkpoint based on the noisy numbers. The best classifier is used to calculate the test set accuracy. Since the training and validation sets are two disjoint sets, according to the parallel composition theorem (McSherry, 2009), the ultimate privacy guarantee depends only on the worst of the guarantees of each analysis, , where is the privacy costs of the validation set and is the privacy cost of training the synthesizer on the training set. DPImageBench sets without additional privacy cost. Thus, the final privacy cost is still (Dwork et al., 2014). We abbreviate this method as ‘Noisy SenV’.
DPImageBench provides ‘SynV’ and ‘Noisy SenV’, the two methods that satisfy DP, avoid bias, and support tuning, for evaluation.
5.3.5. Empirical DP Testing
Empirical DP testing experimentally audits an algorithm’s privacy by running it on specially constructed data to assess DP parameters and measure potential exposure of personal information. The goal is to determine whether the system effectively obscures sensitive details and adheres to DP (Nasr et al., 2021).
We chose not to perform empirical DP testing for the following reasons. (1) Lack of suitable methods: Most existing research focuses on auditing DP in classification models (Nasr et al., 2021; Annamalai and De Cristofaro, 2024), models trained on non-image data (e.g., tabular data (Annamalai et al., 2024)), or the DP-SGD algorithm (Steinke et al., 2023; Jagielski et al., 2020). Applying these methods to our problem is not straightforward, as we focus on DP image generation, and some of the approaches we evaluate, such as PE, do not use DP-SGD. (2) Theoretical guarantee: DP synthesis algorithms already have the rigorous theoretical privacy guarantee that the sensitive information leaked from the output data is bounded. If the goal is empirical privacy, we do not even need DP in the first place.
6. Experimental Setup
Investigated Datasets. Drawing on prior work (Li et al., 2024; Lin et al., 2024; Ghalebikesabi et al., 2023; Wang et al., 2024; Dockhorn et al., 2023), we categorize datasets into sensitive datasets requiring protection and public datasets that can be shared without privacy concerns. Appendix C Table 16 summarizes the datasets, noting that some public datasets like LAION (Schuhmann et al., 2022) raise privacy and copyright issues, making them unsuitable for pretraining (51CTO, 2022).
We select seven widely studied sensitive datasets: (1) MNIST (LeCun et al., 1998), (2) FashionMNIST (Xiao et al., 2017), (3) CIFAR-10 (Krizhevsky, 2009), (4) CIFAR-100 (Krizhevsky, 2009), (5) EuroSAT (Helber et al., 2019), (6) CelebA (Liu et al., 2015), and (7) Camelyon (Bándi et al., 2019), commonly used in prior studies (Li et al., 2024; Lin et al., 2024; Liu et al., 2023). The ImageNet ISLVRC2012 (Deng et al., 2009) and Places365 (Zhou et al., 2017) datasets are treated as public datasets. For CelebA (Liu et al., 2015), we focus on gender classification tasks following(Li et al., 2024). MNIST and FashionMNIST consist of grayscale images, while the others are in color, with further details in Appendix C.
Implementation of Evaluation Metrics. This section describes the metrics for evaluating synthetic images. We generate 60,000 synthetic images per sensitive dataset for evaluation. For metrics like FID, IS, Precision, and Recall, we use the pretrained Inception v3 (Szegedy et al., 2016) from Github to extract features or classify images.222https://github.com/mseitzer/pytorch-fid/releases/download/fid_weights/pt_inception-2015-12-05-6726825d.pth Hyper-parameter settings for utility and fidelity metrics are detailed in Appendices B.2.4 and B.2.5. We assess the utility of synthetic images in image classification tasks using DPImageBench, which employs three classification algorithms—ResNet (He et al., 2016), WideResNet (Zagoruyko, 2016), and ResNeXt (Xie et al., 2017)—trained on synthetic images, with accuracy evaluated on the sensitive test set.
@@ 7. EMPIRICAL EVALUATIONS
We provide the computational resources and runtime analysis of studied methods in Table 18 of Appendix D.2.
7.1. The Quality of Synthetic Images (RQ1)
@@
Experiment Design. To address RQ1, this section evaluates the utility and fidelity of synthetic images generated by the studied methods under privacy budgets , with visualizations for intuitive insights. The synthesizer sizes for GAN-based and diffusion-based methods are 5.4M and 3.8M, respectively. For GAN-based methods, we use BigGAN (Brock et al., 2019) with a generator size of 3.8M and a discriminator size of 1.6M, aligning the generator sizes with diffusion-based methods. For diffusion-based methods, we adopt the model size settings from DPDM (Dockhorn et al., 2023) and PrivImage (Li et al., 2024). In latent diffusion-based methods (e.g., DP-LDM and DP-LoRA), as described in Section 2.2, an autoencoder is integrated into the diffusion process, with DMs sized at 3.8M and the autoencoder following the original implementations in DP-LDM (Liu et al., 2023) and DP-LoRA (Tsai et al., 2024) (see Appendix B.2.2 for network architecture details).
@@
High-dimensional sensitive datasets with resolutions above are resized to for fair comparison, following prior work (Ghalebikesabi et al., 2023; Li et al., 2024; Dockhorn et al., 2023). For CelebA, we also tested higher resolutions ( and ). Unless specified, utility evaluations in this paper use the ‘Noisy SenV’ method.
@@
Result Analysis. Tables 5 and 6 present the Acc of classifiers trained on synthetic images, with the best score per column highlighted in a gray box. Table 7 reports fidelity results under . Note that our experimental settings differ from the original papers in model architecture, failure probability , and evaluation methods, which may cause performance variations compared to the reported results. Figure 9 in Appendix D.3 provides examples of synthetic images. These findings reveal the following key takeaways:
@@
Directly tuning classifiers using the sensitive dataset overestimates the utility of synthetic images. For instance, in Table 5, DP-Kernel on EuroSAT shows an accuracy drop of 21.2% and 24.9% under when using the synthetic dataset as the validation set. Tables 5 and 6 present the average results for Noisy SenV’ and SynV,’ showing a decrease of 1.33% and 4.48% in accuracy under , and 1.12% and 3.00% under , compared to the ‘SenV’ evaluation method. As increases, the quality of synthetic images improves, leading to reduced utility evaluation errors. These results also indicate that the ‘Noisy Senv’ method achieves better utility evaluation than the ‘SynV.’
@@
Our fidelity and utility evaluation provides a more comprehensive view of the quality of synthetic images. IS evaluates the practical quality of synthetic images but struggles with unnatural datasets like MNIST, FashionMNIST, and EuroSAT (Jiralerspong et al., 2023). In Table 7, PE achieves IS scores of 2.79, 5.37, 14.97, 19.65, 8.16, 3.87, and 4.58 for the seven sensitive datasets; six out of seven datasets achieve the highest IS across all methods. However, despite the high practical quality of these images, they are not of good utility, achieving Acc scores of 32.7%, 57.8%, 75.3%, 24.9%, 36.8%, 74.2%, and 64.9%. Five out of seven of them are lower than the best methods.
For DP-LDM and DP-LoRA for MNIST, the Acc under is 95.5% and 97.1%, but the FID is 99.1 and 95.4. The synthetic images retain essential information for downstream models but may lack realistic textures, colors, and fine details, as these methods prioritize class-discriminative features like shapes over photorealism. In their original implementations, DP-LDM and DP-LoRA resize MNIST images to , whereas the original size of MNIST image is . This resizing results in better FID scores reported in their paper compared to DPImageBench. When resizing to , DP-LDM and DP-LoRA achieve FID scores of 13.4 and 12.1, similar to what was reported in their papers. Precision and Recall are decoupled from FID, meaning that having one value high and the other low does not necessarily improve the FID for synthetic images. For example, in the case of DPGAN with CIFAR-10, the precision is 0.67, the best among all methods. However, its recall is as low as 0.01, and the FID is as high as 138.7. FLD considers the novelty of synthetic images, and better FLD scores generally indicate that most other fidelity metrics are also favorable. The human perspective metric, IR, is consistently negative due to the low resolution of synthetic images compared to in paper (Xu et al., 2024). Further fidelity evaluation details for Camelyon are in Appendix D.1.
@@
Current methods still face challenges in high-resolution DP image synthesis. Table 8 shows the Acc and FID of synthetic CelebA with varying resolutions. As resolutions of synthetic images increases from to , average accuracy and FID drop from 84.0 and 87.5 to 65.6 and 218.4, respectively, indicating that current methods struggle with high-resolution synthesis. However, PE is the least affected by the increased resolutions in synthetic images, as the synthetic images of PE only depend on the capability of foundational APIs. Other methods, such as PrivImage and DP-LDM, which train synthesizers using DP-SGD, fail in high-resolution image synthesis. High-resolution images contain more complex spatial structures, requiring more precise gradients to capture meaningful patterns. However, DP-SGD injects noise into gradients, which disrupts these updates. As image resolution increases, the number of model parameters and/or the number of required model updates also grows, making it harder for DP-SGD to converge. We use a default synthesizer size of 3.8M; larger sizes may enhance performance.
7.2. Model Size and Privacy Budget (RQ2)